The investment committee of Sentry Insurance Co. is evaluating two projects, office expansion and upgrade to computer servers. The projects have different useful lives, but each requires an investment of $490,000. The estimated net cash flows from each project are as follows:

Net Cash Flows
Year Office Expansion Servers
1 $125,000 $165,000
2   125,000   165,000
3   125,000   165,000
4   125,000   165,000
5   125,000
6   125,000

The committee has selected a rate of 12% for purposes of net present value analysis. It also estimates that the residual value at the end of each project's useful life is $0, but at the end of the fourth year, the office expansion's residual value would be $180,000.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

1. For each project, compute the net present value. Use the present value of an annuity of $1 table above. Ignore the unequal lives of the projects. If required, round to the nearest dollar.

Office Expansion Server Upgrade
Present value of annual net cash flows $ $
Less amount to be invested
Net present value $ $

2. For each project, compute the net present value, assuming that the office expansion is adjusted to a four-year life for purposes of analysis. Use the present value of $1 table above.

Office Expansion Server Upgrade
Present value of net cash flow total $ $
Less amount to be invested
Net present value $ $
Answer #1
Office Expansion Server Upgrade
Present value of annual net cash flows               513,875               501,105
Less amount to be invested               490,000               490,000
Net present value                 23,875                 11,105
2.Four year life
Office Expansion Server Upgrade
Present value of annual net cash flows               494,105               501,105
Less amount to be invested               490,000               490,000
Net present value                   4,105                 11,105

B Office Expansion C Server Upgrade 1 2 Present value of annual net cash flows 1=125000*120 |=165000-118 3 Less amount to be